OpenStudy (anonymous):

8.9 = 2/3 log 10 of (K) - 0.9 K = energy released, how do I work this out please :)

OpenStudy (anonymous):

So this is just algebra. Add .9 on both sides. 9.8 =2/3 log (base 10) of k Then you multiply both sides by 3/2 14.7 = log(base 10) of k A log of base 10 means you use 10 as a base for an exponential number Thusly, 10^ 14.7 = k So 10 to the power of 14.7 is equal to k.

OpenStudy (anonymous):

Thank you so very much.....I was there until the last step and wasn't sure what to do when you put the log to the other side, thanks so much again :)